OpenStudy (anonymous):

squareroot of 9 / squareroot of 3

OpenStudy (hoblos):

sqrt9 = 3 sqrt9/sqrt3 = 3/sqrt3 = sqrt3

OpenStudy (anonymous):

\[\sqrt{9}\div \sqrt{3}\]

OpenStudy (hoblos):

\[\sqrt{9}/\sqrt{3}= 3/\sqrt{3} = \sqrt{3}\]

OpenStudy (anonymous):